OpenStudy (anonymous):

Please evaluate the following integrals. ∫ dx/e^(x-1)

OpenStudy (anonymous):

\[\int\frac{dx}{e^{x-1}}=\int e^{1-x}~dx\] Substitute \(u=1-x\), so that \(-du=dx\): \[-\int e^{u}~du\]

OpenStudy (anonymous):

What is the answer sir?

OpenStudy (anonymous):

That's for you to find out. I've given enough hints.

OpenStudy (anonymous):

another hint please. Thank you so much

OpenStudy (primeralph):

I solved this already for you.

OpenStudy (anonymous):

I didnt get it sir. ;(

OpenStudy (primeralph):